OpenStudy (anonymous):

Which of the following is the radical expression of 4 times d to the three eighths power?

OpenStudy (anonymous):

\[\sqrt[8]{4d^3}\]

OpenStudy (anonymous):

really depends on what is being raised to the power of \(\frac{3}{8}\) the \(4d\) or just the \(d\)

OpenStudy (anonymous):

good point @satellite73 i didnt even realize that

OpenStudy (anonymous):

if it is the \(4d\) then your answer is \[\sqrt[8]{(4d)^3}\]

OpenStudy (anonymous):

if it is just the \(d\) i.e. you see \[4d^{\frac{3}{8}}\] then it will be \[4\sqrt[8]{d^3}\]